OpenStudy (rosales0228):

i have no clue how to do this at all!:/ Solve this equation for x. Round your answer to the nearest hundredth. 0.69=log x

OpenStudy (rosales0228):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

logs undo exponents and vice versa

jimthompson5910 (jim_thompson5910):

so something like log(x) = 2 turns into x = 10^2 where the log above is by default base 10 (if no base is stated otherwise)

jimthompson5910 (jim_thompson5910):

assuming you're dealing with base 10, 0.69=log x will transform into x = 10^(0.69) x = ??